# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ 4 Opening lenses
Let the function $g(z)$ be the unique solution to RHP 3.3. For $z\in\eta_{1}$, we have the representation
$g_{+}(z)-g_{-}(z)=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R_{-}(s)}{\rm d}s.$ (4.1)
The quantity $g_{+}-g_{-}$ has an analytic extension to $\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right)$. We define $u(z)$ to be that analytic extension,
$u(z)=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R(s)}{\rm d}s,\ \qquad z\in\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right),$ (4.2)
where the path of integration is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$, so that
$u_{-}(z)=g_{+}(z)-g_{-}(z),\ \ z\in\eta_{1}.$ (4.3)
There exists a contour $\eta_{3}$ connecting $A$ with $-{\overline{A}}$ such that
$u(z)\in i\mathbb{R},\qquad z\in\eta_{3}.$ (4.4)
The contour $\eta_{3}$ lies above $\eta_{1}$ and we take it to be oriented from $A$ to $-{\overline{A}}$ (like $\eta_{1}$).
Similarly, for $z\in\eta_{2}$, we have
$g_{+}(z)-g_{-}(z)=\frac{4c}{\tau}\int_{A}^{-\overline{A}}\frac{1}{R_{-}(s)}{\rm d}s+\frac{4c}{\tau}\int_{-A}^{z}\frac{1}{R_{-}(s)}{\rm d}s.$ (4.5)
@@
where again the path is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Note that
$\tilde{u}_{-}(z)=\frac{4c}{\tau}\int_{A}^{-\overline{A}}\frac{1}{R_{-}(s)}{\rm d}s+\frac{4c}{\tau}\int_{-A}^{z}\frac{1}{R_{-}(s)}{\rm d}s,\ \qquad z\in\eta_{2},$ (4.7)
so that
$\tilde{u}_{-}(z)=g_{+}(z)-g_{-}(z),\ \ z\in\eta_{2}.$ (4.8)
There exists a contour $\eta_{4}$ connecting $-A$ with ${\overline{A}}$ such that
$\tilde{u}(z)\in i\mathbb{R},\qquad z\in\eta_{4}.$ (4.9)
The contour $\eta_{4}$ lies below $\eta_{2}$ and we take it to be oriented from $-A$ to ${\overline{A}}$ (like $\eta_{2}$).
Furthermore, the following hold
$\frac{2c}{\tau}\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s=1,\quad\frac{2c}{\tau}\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s=\frac{1}{\tau},\quad-\frac{2c}{\tau}\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s=\frac{1}{\tau}-1\,.$ (4.10)
@@
    $J_{D}(z)=\begin{cases}\begin{pmatrix}e^{-\ln(N)u(z)+\widetilde{y}_{-}(z)-\widetilde{y}_{+}(z)}&0\\ \\[2.0pt] i&e^{\ln(N)u(z)+\widetilde{y}_{+}(z)-\widetilde{y}_{-}(z)}\end{pmatrix},&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}e^{-\ln(N)\tilde{u}(z)+\widetilde{y}_{-}(z)-\widetilde{y}_{+}(z)}&i\\ \\[2.0pt] 0&e^{\ln(N)\tilde{u}(z)+\widetilde{y}_{+}(z)-\widetilde{y}_{-}(z)}\end{pmatrix},&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta}\end{pmatrix},&z\in\gamma\end{cases}$
@@
$\gamma$$\eta_{3}$$\eta_{4}$$\eta_{1}$$\eta_{2}$$\Omega_{3}^{+}$$\Omega_{2}^{+}$
@@ 4.1 Jump matrix factorization
$J_{D}(z)=\begin{cases}\begin{pmatrix}1&-ie^{-\ln(N)u(z)}\frac{e^{2\widetilde{y}_{-}(z)}}{2\pi h(z)\rho(z)e^{2i\theta(z)}}\\ 0&1\end{pmatrix}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\begin{pmatrix}1&-ie^{\ln(N)u(z)}\frac{e^{2\widetilde{y}_{+}(z)}}{2\pi h(z)\rho(z)e^{2i\theta(z)}}\\ 0&1\end{pmatrix},\quad&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\tilde{u}(z)}\frac{e^{-2\widetilde{y}_{-}(z)}}{2\pi h(z)\rho(z)e^{-2i\theta(z)}}&1\end{pmatrix}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\begin{pmatrix}1&0\\ -ie^{-\ln(N)\tilde{u}(z)}\frac{e^{-2\widetilde{y}_{+}(z)}}{2\pi h(z)\rho(z)e^{-2i\theta(z)}}&1\end{pmatrix},\quad&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta}\end{pmatrix},\quad&z\in\gamma\end{cases}$
@@
$J_{D}(z)=\begin{cases}\left(\mathfrak{M}^{-}\right)_{-}(z)\begin{pmatrix}0&i\\ i&0\end{pmatrix}\left(\mathfrak{M}^{+}\right)_{+}(z),\quad&z\in\eta_{3}\\ \\[2.0pt] \left(\mathfrak{N}^{+}\right)_{-}(z)\begin{pmatrix}0&i\\ i&0\end{pmatrix}\left(\mathfrak{N}^{-}\right)_{+}(z),\quad&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta}\end{pmatrix},\quad&z\in\gamma\end{cases}$
@@
    $J_{E}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{3}\cup\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta}\end{pmatrix},\quad&z\in\gamma\\ \\[4.0pt] \mathfrak{M}^{-}(z),\quad&z\in\omega_{3}\cup\omega_{4}\\ \\[4.0pt] \mathfrak{N}^{+}(z),\quad&z\in\omega_{1}\cup\omega_{2}\end{cases}$
@@
    $E(z)=I_{2}+\mathcal{O}(1/z)$, as $z\to\infty$.
$\gamma$$\eta_{3}$$\eta_{4}$$\omega_{1}$$\omega_{2}$$\omega_{3}$$\omega_{4}$$\Omega_{1}$$\Omega_{2}^{-}$$\Omega_{3}^{-}$$\Omega_{4}$$\eta_{1}$$\eta_{2}$$\Omega_{3}^{+}$$\Omega_{2}^{+}$
@@ 5 Global Parametrix
    $J_{\Lambda}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\,,&z\in\eta_{3}\cup\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{-\Delta+\frac{2\ln(N)}{\tau}}&0\\ 0&e^{\Delta-\frac{2\ln(N)}{\tau}}\end{pmatrix},&z\in\gamma\end{cases}$ (5.1)
@@
    $\Lambda(z)$ satisfies the asymptotic behavior: $\Lambda(z)=I_{2}+\mathcal{O}(1/z)$, as $z\to\infty$
@@
    $\nu_{+}(z)-\nu_{-}(z)=i\left(\Delta-2\frac{\ln(N)}{\tau}\right),\quad z\in\gamma$.
@@
$\phi_{-}(z)-\phi_{+}(z)=\frac{i\left(\Delta-2\frac{\ln(N)}{\tau}\right)}{S(z)},\quad z\in\gamma$ (5.4)
@@
$\phi(z)=\frac{1}{2\pi}\int_{\gamma}\frac{1}{S(s)}\frac{\Delta-2\frac{\ln(N)}{\tau}}{s-z}\,{\rm d}s.$ (5.5)
@@
$\nu(z)=\frac{S(z)}{2\pi}\int_{\gamma}\frac{1}{S(s)}\frac{\Delta-2\frac{\ln(N)}{\tau}}{s-z}\,{\rm d}s.$ (5.6)
@@
$a_{1}=-\frac{1}{2\pi}\int_{\gamma}\frac{\Delta-\frac{2\ln(N)}{\tau}}{S(s)}\,{\rm d}s=-\frac{\tau\Delta-2\ln(N)}{4c\pi},\quad a_{0}=-\frac{1}{2\pi}\int_{\gamma}\frac{\Delta-\frac{2\ln(N)}{\tau}}{S(s)}\,s\,{\rm d}s.$ (5.8)
@@
    $\Xi(z)=\left[I_{2}+\mathcal{O}(1/z)\right]e^{-i(a_{1}z+a_{0}+\nu_{0})\sigma_{3}},\quad z\to\infty$
@@
$W(p)=\left[\begin{pmatrix}1\\ 0\end{pmatrix}+\mathcal{O}(1/p)\right]e^{-i(a_{1}p+a_{0}+\nu_{0})},\quad p\to\infty_{1}$ (5.12a)
$W(p)=\left[\begin{pmatrix}0\\ 1\end{pmatrix}+\mathcal{O}(1/p)\right]e^{i(a_{1}p+a_{0}+\nu_{0})},\quad p\to\infty_{2}$ (5.12b)
where $\infty_{1}$ and $\infty_{2}$ are the images of $\infty$ in $S_{I}$ and $S_{II}$, respectively. Following [6], we solve RHP 5.4 in terms of Theta functions. To this end, we define the function $\rho$ as follows
@@
and we recall that it is positive imaginary, see Remark 3.1. Next, we introduce the Abel map related to ${\mathcal{R}}$ as follows
$\mathcal{A}(p)=\int_{A}^{p}\frac{c}{\rho(s)}\,{\rm d}s.$ (5.15)
This is typically multi-valued, and if one wants to create a single-valued function, one may restrict the path of integration to avoid $P_{I}(\gamma)\cup P_{II}(\gamma)\cup P_{I}([{\overline{A}},\infty])\cup P_{II}([{\overline{A}},\infty])$. Moreover, the function $\mathcal{A}$ satisfies the following conditions
$\mathcal{A}_{+}(p)+\mathcal{A}_{-}(p)=0,\quad p\in\eta_{3},$ (5.16)
$\mathcal{A}_{+}(p)-\mathcal{A}_{-}(p)=-1,\quad p\in P_{I}(\gamma),$ (5.17)
$\mathcal{A}_{+}(p)-\mathcal{A}_{-}(p)=1,\quad p\in P_{II}(\gamma),$ (5.18)
$\mathcal{A}_{+}(p)+\mathcal{A}_{-}(p)=\tau,\quad p\in\eta_{4}$ (5.19)
Next, we consider the Jacobi Theta-$3$ function [12, 20.2]
which is an even function of $z$ satisfying the following properties
$\Theta(z+1;\tau)=\Theta(z;\tau)$ (5.21)
$\Theta(z+\tau;\tau)=e^{-\pi i\tau-2\pi iz}\Theta(z;\tau)$ (5.22)
and has a simple zero at $z=\frac{\tau+1}{2}$. We use the $\Theta$-function to construct $W(p)$ for $p\in{\mathcal{R}}$, and then through (5.11) we find $\Xi(z)$ solving RHP 5.4. Recall that the function $W(p)$ has simple poles at the points ${\overline{A}},-{\overline{A}}$ and that two entries of $W(p)$ vanish at $\infty_{1}$ and $\infty_{2}$.
We first construct two meromorphic scalar functions on ${\mathcal{R}}$, each having simple poles at ${\overline{A}},-{\overline{A}}$, with one function vanishing at $\infty_{1}$ and the other at $\infty_{2}$. We guess that such functions have the following form
$\begin{split}&F_{1}(p;\tau)=\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{2})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{1})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}\\ &F_{2}(p;\tau)=\frac{\Theta\left(\mathcal{A}(p)-\mathcal{A}(\infty_{1})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(Q_{2})+\frac{\tau+1}{2};\tau\right)}{\Theta\left(\mathcal{A}(p)-\mathcal{A}({\overline{A}})+\frac{\tau+1}{2};\tau\right)\Theta\left(\mathcal{A}(p)-\mathcal{A}(-{\overline{A}})+\frac{\tau+1}{2};\tau\right)}\end{split}$ (5.24)
where $Q_{1},Q_{2}\in{\mathcal{R}}$ are zeros of $F_{1}(p;\tau)$ and $F_{2}(p;\tau)$, respectively, on ${\mathcal{R}}$, which will be determined. The structure of $F_{1}$ and $F_{2}$ ensures that the number of poles of each matches the number of zeros (which must be true for meromorphic functions in this Riemann surface). Additionally, they do not have jumps across $\eta_{3},\eta_{4}$, but they do across $\gamma$. Thus, they nearly solve the RHP for $W(p)$, except for the asymptotic behavior of $F_{1}$ at $\infty_{1}$, $F_{2}$ at $\infty_{2}$, and their jump across $\gamma$. To fix the asymptotic behavior of $W(p)$, we introduce the function $\delta(p)$ on ${\mathcal{R}}$ as follows
@@
$\begin{split}{\widetilde{a}}_{0}=-2ca_{1}\int_{A}^{-{\overline{A}}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\,,\quad{\widetilde{a}}_{1}=0\,,\quad{\widetilde{a}}_{2}=a_{1}\,,\\ \nu_{0}=a_{1}(1-\tau)\int_{A}^{-{\overline{A}}}\frac{s^{2}}{S_{+}(s)}{\rm d}s-a_{1}A-a_{0}+a_{1}\int_{A}^{\infty}\frac{s^{2}-S(s)}{S(s)}{\rm d}s\,,\end{split}$ (5.26)
@@
$\delta_{+}(z)+\delta_{-}(z)=0,\quad z\in\eta_{3}$ (5.28)
$\delta_{+}(z)+\delta_{-}(z)=\zeta(x,N),\quad z\in\eta_{4}$ (5.29)
@@
$\zeta(x,N)=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right)\in i{\mathbb{R}}.$ (5.30)
@@
${\widetilde{a}}_{2}=a_{1}\,,\qquad{\widetilde{a}}_{1}=0\,,\qquad\nu_{0}=\frac{\tau-1}{2c}{\widetilde{a}}_{0}-a_{1}A-a_{0}+a_{1}\int_{A}^{\infty}\frac{s^{2}-S(s)}{S(s)}{\rm d}s\,,$ (5.32)
@@
$\begin{split}\delta_{+}(p)+\delta_{-}(p)&=2a_{1}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{S(s)}{\rm d}s-\tau\int_{\eta_{3}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\right)\\ &=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{S(s)}{\rm d}s-\tau\int_{\eta_{3}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\right)\\ &=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right)=:\zeta(x,N)\,.\end{split}$ (5.38)
We used Lemma 4.1 to simplify the expression. We notice that $\zeta(x,N)$ depends on $N$ and $x$ since $a_{1}$ does. Furthermore, it is easy to check that $\zeta\in i{\mathbb{R}}$. This concludes the proof of the proposition. ∎

### v2

@@ 4 Opening lenses
Let the function $g(z)$ be the unique solution to RHP 3.3. Define the function $u(z)$ for $z\in{\mathbb{C}}_{+}\setminus(\eta_{1}\cup\gamma)$ as
$u(z)=-2g(z)+1=\frac{4c}{\tau}\int_{A}^{z}\frac{1}{R(s)}{\rm d}s,\ \qquad z\in\mathbb{C}_{+}\setminus\left(\eta_{1}\cup\gamma\right),$ (4.1)
where the path of integration is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Let $\eta_{3}$ be the arc of the circle of radius $|A|$ centered at 00 connecting $A$ and $-\overline{A}$. We have
$u_{-}(z)=g_{+}(z)-g_{-}(z),\quad z\in\eta_{1}\,,\qquad u(z)\in i\mathbb{R},\quad z\in\eta_{3}$ (4.2)
@@
where again the path is chosen to avoid $\eta_{1}\cup\eta_{2}\cup\gamma$. Let $\eta_{4}$ be the arc of the circle of radius $|A|$ centered at 00 connecting $-A$ and $\overline{A}$. We have
${\widetilde{u}}_{-}(z)=g_{+}(z)-g_{-}(z),\quad z\in\eta_{2}\,,\qquad u(z)\in i\mathbb{R},\quad z\in\eta_{4}$ (4.4)
Clearly $\eta_{4}$ lies below $\eta_{2}$ and we take it to be oriented from $-A$ to ${\overline{A}}$ (like $\eta_{2}$).
@@
    $J_{D}(z)=\begin{cases}\begin{pmatrix}e^{-\ln(N)u(z)+\widetilde{y}_{-}(z)-\widetilde{y}_{+}(z)}&0\\ \\[2.0pt] i&e^{\ln(N)u(z)+\widetilde{y}_{+}(z)-\widetilde{y}_{-}(z)}\end{pmatrix},&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}e^{-\ln(N)\tilde{u}(z)+\widetilde{y}_{-}(z)-\widetilde{y}_{+}(z)}&i\\ \\[2.0pt] 0&e^{\ln(N)\tilde{u}(z)+\widetilde{y}_{+}(z)-\widetilde{y}_{-}(z)}\end{pmatrix},&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta(x)}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta(x)}\end{pmatrix},&z\in\gamma\end{cases}$
@@
$\gamma$$\eta_{1}$$\eta_{2}$$\eta_{3}$$\eta_{4}$$\Omega_{2}^{+}$$\Omega_{3}^{+}$
@@ 4.1 Jump matrix factorization
$J_{D}(z)=\begin{cases}\begin{pmatrix}1&-ie^{-\ln(N)u(z)}\frac{e^{2\widetilde{y}_{-}(z)}}{2\pi h(z)\rho(z)e^{2i\theta(z)}}\\ 0&1\end{pmatrix}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\begin{pmatrix}1&-ie^{\ln(N)u(z)}\frac{e^{2\widetilde{y}_{+}(z)}}{2\pi h(z)\rho(z)e^{2i\theta(z)}}\\ 0&1\end{pmatrix},\quad&z\in\eta_{3}\\ \\[4.0pt] \begin{pmatrix}1&0\\ -ie^{\ln(N)\tilde{u}(z)}\frac{e^{-2\widetilde{y}_{-}(z)}}{2\pi h(z)\rho(z)e^{-2i\theta(z)}}&1\end{pmatrix}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\begin{pmatrix}1&0\\ -ie^{-\ln(N)\tilde{u}(z)}\frac{e^{-2\widetilde{y}_{+}(z)}}{2\pi h(z)\rho(z)e^{-2i\theta(z)}}&1\end{pmatrix},\quad&z\in\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta(x)}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta(x)}\end{pmatrix},\quad&z\in\gamma\end{cases}$
@@
$J_{D}(z)=\begin{cases}\left(\mathfrak{M}^{-}\right)_{-}(z)\begin{pmatrix}0&i\\ i&0\end{pmatrix}\left(\mathfrak{M}^{+}\right)_{+}(z),\quad&z\in\eta_{3}\\ \\[2.0pt] \left(\mathfrak{N}^{+}\right)_{-}(z)\begin{pmatrix}0&i\\ i&0\end{pmatrix}\left(\mathfrak{N}^{-}\right)_{+}(z),\quad&z\in\eta_{4}\\ \\[2.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta(x)}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta(x)}\end{pmatrix},\quad&z\in\gamma\end{cases}$
@@
    $J_{E}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix},\quad&z\in\eta_{3}\cup\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{\frac{2}{\tau}\ln(N)-\Delta(x)}&0\\ 0&e^{-\frac{2}{\tau}\ln(N)+\Delta(x)}\end{pmatrix},\quad&z\in\gamma\\ \\[4.0pt] \mathfrak{M}^{-}(z),\quad&z\in\omega_{3}\cup\omega_{4}\\ \\[4.0pt] \mathfrak{N}^{+}(z),\quad&z\in\omega_{1}\cup\omega_{2}\end{cases}$
@@
    $E(z)=I+\mathcal{O}(1/z)$, as $z\to\infty$.
$\gamma$$\eta_{1}$$\eta_{2}$$\eta_{3}$$\eta_{4}$$\Omega_{2}^{+}$$\Omega_{3}^{+}$$\Omega_{2}^{-}$$\Omega_{1}$$\Omega_{2}$$\Omega_{3}^{-}$$\omega_{1}$$\omega_{2}$$\omega_{3}$$\omega_{4}$
@@ 5 Global Parametrix
    $J_{\Lambda}(z)=\begin{cases}\begin{pmatrix}0&i\\ i&0\end{pmatrix}\,,&z\in\eta_{3}\cup\eta_{4}\\ \\[4.0pt] \begin{pmatrix}e^{-\Delta(x)+\frac{2\ln(N)}{\tau}}&0\\ 0&e^{\Delta(x)-\frac{2\ln(N)}{\tau}}\end{pmatrix},&z\in\gamma\end{cases}$ (5.1)
@@
    $\Lambda(z)$ satisfies the asymptotic behavior: $\Lambda(z)=I+\mathcal{O}(1/z)$, as $z\to\infty$
@@
    $\nu_{+}(z)-\nu_{-}(z)=i\left(\Delta(x)-2\frac{\ln(N)}{\tau}\right),\quad z\in\gamma$.
@@
$\phi_{-}(z)-\phi_{+}(z)=\frac{i\left(\Delta(x)-2\frac{\ln(N)}{\tau}\right)}{S(z)},\quad z\in\gamma$ (5.4)
@@
$\phi(z)=\frac{1}{2\pi}\int_{\gamma}\frac{1}{S(s)}\frac{\Delta(x)-2\frac{\ln(N)}{\tau}}{s-z}\,{\rm d}s.$ (5.5)
@@
$\nu(z)=\frac{S(z)}{2\pi}\int_{\gamma}\frac{1}{S(s)}\frac{\Delta(x)-2\frac{\ln(N)}{\tau}}{s-z}\,{\rm d}s.$ (5.6)
@@
$a_{1}=-\frac{1}{2\pi}\int_{\gamma}\frac{\Delta(x)-\frac{2\ln(N)}{\tau}}{S(s)}\,{\rm d}s=-\frac{\tau\Delta(x)-2\ln(N)}{4c\pi},\quad a_{0}=-\frac{1}{2\pi}\int_{\gamma}\frac{\Delta(x)-\frac{2\ln(N)}{\tau}}{S(s)}\,s\,{\rm d}s.$ (5.8)
@@
    $\Xi(z)=\left[I+\mathcal{O}(1/z)\right]e^{-i(a_{1}z+a_{0}+\nu_{0})\sigma_{3}},\quad z\to\infty$
@@
$W(p)=\left[\begin{pmatrix}1\\ 0\end{pmatrix}+\mathcal{O}(1/p)\right]e^{-i(a_{1}p+a_{0}+\nu_{0})},\quad p\to\infty_{1}$ (5.12)
$W(p)=\left[\begin{pmatrix}0\\ 1\end{pmatrix}+\mathcal{O}(1/p)\right]e^{i(a_{1}p+a_{0}+\nu_{0})},\quad p\to\infty_{2}$ (5.13)
where $\infty_{1}$ and $\infty_{2}$ are the images of $\infty$ in $S_{I}$ and $S_{II}$, respectively. Following [7], we solve RHP 5.4 in terms of Theta functions.
To this end, we define the function $\rho$ as follows
@@
and we recall that it is positive imaginary, see Remark 3.1. Next, we fix the asymptotic behavior of $W(p)$ by introducing the function $\delta(p)$ on ${\mathcal{R}}$ as
@@
$\begin{split}{\widetilde{a}}_{0}=-2ca_{1}\int_{A}^{-{\overline{A}}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\,,\quad{\widetilde{a}}_{1}=0\,,\quad{\widetilde{a}}_{2}=a_{1}\,,\\ \nu_{0}=\frac{a_{1}}{2}(1-\tau)\int_{A}^{-{\overline{A}}}\frac{s^{2}}{S_{+}(s)}{\rm d}s-a_{1}A-a_{0}+a_{1}\int_{A}^{\infty}\frac{s^{2}-S(s)}{S(s)}{\rm d}s\,,\end{split}$ (5.17)
@@
$\delta_{+}(z)+\delta_{-}(z)=0,$ $z\in\eta_{3}$ (5.19)
$\delta_{+}(z)+\delta_{-}(z)=\zeta(x,N),$ $z\in\eta_{4}$ (5.20)
@@
$\zeta(x,N)=i\left(4c\pi x+2\ln(N)+4c\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\right)\in i{\mathbb{R}}.$ (5.21)
@@
${\widetilde{a}}_{2}=a_{1}\,,\qquad{\widetilde{a}}_{1}=0\,,\qquad\nu_{0}=\frac{\tau-1}{4c}{\widetilde{a}}_{0}-a_{1}A-a_{0}+a_{1}\int_{A}^{\infty}\frac{s^{2}-S(s)}{S(s)}{\rm d}s\,,$ (5.23)
@@
$\begin{split}\delta_{+}(p)+\delta_{-}(p)&=2a_{1}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{S(s)}{\rm d}s-\tau\int_{\eta_{3}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\right)\\ &=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{S(s)}{\rm d}s-\tau\int_{\eta_{3}}\frac{s^{2}}{S_{+}(s)}{\rm d}s\right)\\ &=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right)\,.\end{split}$ (5.29)
We used Lemma 4.1 to simplify the expression. Applying the Riemann bilinear relation [30] one can rewrite the previous jumps as
$\begin{split}\delta_{+}(p)+\delta_{-}(p)&:=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right)\\ &=i\left(4c\pi x+2\ln(N)+4c\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\right)\,=\zeta(x,N)\,,\end{split}$ (5.30)
from which it follows that $\zeta\in i{\mathbb{R}}$. This concludes the proof of the proposition.